### 06 Mar 2026 v1

# Haar-Type Measures on Topological Quasigroups and Kunen’s Theorem

Abstract
Haar measure is a fundamental structure in harmonic analysis on locally compact groups. Its existence reflects the compatibility between topology and the associative algebraic structure of groups. In this paper we propose a framework for Haar-type measures on topological quasigroups. Since associativity is absent, strict translation invariance is generally too strong to expect. We therefore introduce quasi-invariant measures whose defect is measured by a modular cocycle attached to translations.
We then explain, in a detailed and cautious form, how Moufang-type identities may impose strong constraints on this cocycle. In particular, under additional quasi-invariance assumptions for right translations, the Moufang-type identity $(N1)$ leads naturally to a multiplicativity relation for the cocycle. This suggests a measure-theoretic interpretation of Kunen’s theorem: the emergence of loop structure may be viewed as the collapse of a modular defect in the translation geometry of a quasigroup.
Keywords: topological quasigroup, Haar-type measure, quasi-invariant measure, modular cocycle, Moufang identity, loop, Kunen’s theorem
MSC2020: 20N05, 22A05, 28C10, 43A05
1 Introduction
Haar measure is one of the central structures in the theory of locally compact groups. If $G$ is a locally compact group, then there exists a nonzero regular Borel measure $\mu$ such that
$\mu(gE)=\mu(E)\qquad(g\in G),$
for every Borel set $E\subseteq G$. Equivalently, if $L_{g}:G\to G$ denotes left translation,
$L_{g}(x)=gx,$
then
$(L_{g})_{*}\mu=\mu\qquad(g\in G).$
This compatibility between the topology and the translation structure is one of the foundations of abstract harmonic analysis [3, 4, 5].
A natural question is whether an analogous structure can exist for topological quasigroups. Quasigroups retain the bijectivity of left and right translations, but they lack associativity [8, 9, 10]. Consequently, the classical proof of the Haar theorem cannot be transferred directly. In particular, the family of left translations is no longer modeled by the quasigroup itself in the way it is for groups.
The purpose of this paper is not to claim a full Haar theorem for arbitrary topological quasigroups. Rather, the aim is to formulate a careful measure-theoretic framework in which the failure of strict translation invariance is encoded by a positive cocycle. This cocycle plays the role of a modular defect. We then explain how Moufang-type identities may constrain this defect.
The central motivating example is Kunen’s theorem [2], which states that a quasigroup satisfying the Moufang-type identity $(N1)$ is necessarily a loop. The identity is
$((xy)z)y=x(y(zy)).$
Our guiding idea is that, under suitable quasi-invariance hypotheses, $(N1)$ forces strong compatibility relations among translation operators, and these in turn constrain the modular cocycle. In this sense, loop structure may be viewed as a kind of unimodularity phenomenon in the translation geometry of a quasigroup.
The purpose of this paper is to propose a framework in which Haar-type structures can be studied in the quasigroup setting using quasi-invariant measures and modular cocycles.
This paper proposes a measure-theoretic viewpoint on quasigroup structure, suggesting that loop identities may be interpreted as a form of unimodularity in the translation geometry of topological quasigroups.
Remark 1.
The present paper is primarily conceptual in nature. We do not claim a general existence theorem for Haar-type measures on arbitrary topological quasigroups. Instead, the goal is to isolate a structural mechanism linking translation operators, modular cocycles, and Moufang-type identities.
2 Topological Quasigroups and Translation Operators
Definition 1.
A quasigroup is a set $Q$ equipped with a binary operation
$Q\times Q\to Q,\qquad(x,y)\mapsto xy,$
such that for every $a,b\in Q$ there exist unique elements $x,y\in Q$ with
$ax=b,\qquad ya=b.$
Equivalently, the equations $au=b$ and $va=b$ are uniquely solvable in $u$ and $v$.
Definition 2.
A topological quasigroup is a quasigroup $Q$ endowed with a Hausdorff topology such that multiplication and both division operations are continuous.
For each $a\in Q$ we define the left and right translations
$L_{a}(x)=ax,\qquad R_{a}(x)=xa.$
By the quasigroup axioms, each $L_{a}$ and each $R_{a}$ is a bijection. In a topological quasigroup they are in fact homeomorphisms, since the inverse maps are given by the corresponding division operations.
Remark 2.
For groups one has
$L_{a}\circ L_{b}=L_{ab}.$
This identity is one of the key reasons why Haar measure is available in the group setting. In a quasigroup this relation generally fails, because
$(L_{a}\circ L_{b})(x)=a(bx)$
need not be equal to
$(ab)x=L_{ab}(x).$
Thus the loss of associativity appears immediately at the level of translation operators.
Nevertheless, left and right translations still generate genuine groups of homeomorphisms.
Definition 3.
The left multiplication group of a quasigroup $Q$ is the subgroup
$\mathrm{LMlt}(Q)\leq\mathrm{Homeo}(Q)$
generated by all left translations $L_{a}$ $(a\in Q)$. Similarly, the right multiplication group $\mathrm{RMlt}(Q)$ is the subgroup generated by all right translations, and the multiplication group $\mathrm{Mlt}(Q)$ is generated by both left and right translations.
3 The Moufang Identity and Translation Calculus
We now turn to the Moufang-type identity
$((xy)z)y=x(y(zy)).$
The key point is that this identity can be rewritten as an operator identity on $Q$.
Theorem 1 (Translation form of the Moufang identity).
Let $Q$ be a quasigroup. Then the Moufang identity
$((xy)z)y=x(y(zy))$
holds for all $x,y,z\in Q$ if and only if the translation operators satisfy
$R_{y}\circ L_{xy}=L_{x}\circ L_{y}\circ R_{y}\qquad(x,y\in Q).$
Proof.
Assume first that the Moufang identity
$((xy)z)y=x(y(zy))$
holds for all $x,y,z\in Q$.
Let $x,y\in Q$ be fixed and let $z\in Q$ be arbitrary. We compute the action of the operators on $z$.
First,
$(R_{y}\circ L_{xy})(z)=R_{y}((xy)z)=((xy)z)y.$
On the other hand,
$(L_{x}\circ L_{y}\circ R_{y})(z)=L_{x}(L_{y}(R_{y}(z))).$
Since
$R_{y}(z)=zy,$
we obtain
$L_{y}(R_{y}(z))=y(zy).$
Applying $L_{x}$ gives
$L_{x}(y(zy))=x(y(zy)).$
Therefore
$(L_{x}\circ L_{y}\circ R_{y})(z)=x(y(zy)).$
By the Moufang identity these two expressions coincide, hence
$(R_{y}\circ L_{xy})(z)=(L_{x}\circ L_{y}\circ R_{y})(z)$
for every $z\in Q$.
Thus
$R_{y}\circ L_{xy}=L_{x}\circ L_{y}\circ R_{y}.$
Conversely, suppose the translation identity
$R_{y}\circ L_{xy}=L_{x}\circ L_{y}\circ R_{y}$
holds for all $x,y\in Q$. Applying both sides to an arbitrary element $z\in Q$ gives
$(R_{y}\circ L_{xy})(z)=(L_{x}\circ L_{y}\circ R_{y})(z).$
Expanding both sides yields
$((xy)z)y=x(y(zy)).$
Thus the Moufang identity holds.
4 Haar-Type Measures and Modular Cocycles
4.1 Quasi-invariance under left translations
The natural analogue of Haar measure in the quasigroup setting should interact with the translation structure. Since strict invariance is too strong to expect in general, we begin with quasi-invariance. Let $T:X\to X$ be a measurable map and let $\mu$ be a measure on $X$. The push-forward measure $T_{*}\mu$ is defined by
$(T_{*}\mu)(E)=\mu(T^{-1}(E))$
for every measurable set $E$.
Equivalently, for every nonnegative measurable function $f$,
$\int f(x)\,d(T_{*}\mu)(x)=\int f(T(x))\,d\mu(x).$
Definition 4.
Let $Q$ be a locally compact Hausdorff topological quasigroup. A regular Borel measure $\mu$ on $Q$ is called a left quasi-invariant measure if for every $a\in Q$ there exists a positive scalar $j(a)$ such that
$(L_{a})_{*}\mu=j(a)\mu.$
The function
$j:Q\to\mathbb{R}_{>0}$
is called the left modular cocycle associated with $\mu$.
Remark 3.
The existence of such quasi-invariant measures is a nontrivial problem in general. Unlike the group case, there is currently no general analogue of the Haar existence theorem for topological quasigroups. In this paper we work conditionally under the assumption that such measures are given.
Here $(L_{a})_{*}\mu$ is the push-forward measure defined by
$((L_{a})_{*}\mu)(E)=\mu(L_{a}^{-1}(E))$
for Borel sets $E\subseteq Q$.
The same definition can be expressed at the level of integration. Namely, for every nonnegative Borel function $f$ one has
$\int_{Q}f(x)\,d((L_{a})_{*}\mu)(x)=\int_{Q}f(L_{a}(x))\,d\mu(x).$
Hence the condition
$(L_{a})_{*}\mu=j(a)\mu$
is equivalent to
$\int_{Q}f(ax)\,d\mu(x)=j(a)\int_{Q}f(x)\,d\mu(x)$
for all nonnegative Borel $f$ (and therefore also for all integrable $f$).
Remark 4.
In the group case, if one works with a left Haar measure, then left translations preserve the measure exactly, so the corresponding left cocycle is identically equal to $1$. The nontrivial modular function arises from right translations. In the quasigroup setting we do not have that asymmetry a priori, so it is natural to allow a cocycle already on the left.
4.2 Two-sided quasi-invariance
The basic calculation involving the Moufang identity naturally mixes left and right translations. For that reason, if one wants to perform actual measure-theoretic derivations, it is convenient to impose quasi-invariance for right translations as well.
Definition 5.
A regular Borel measure $\mu$ on a locally compact Hausdorff topological quasigroup $Q$ is called two-sided quasi-invariant if there exist positive functions
$j,\rho:Q\to\mathbb{R}_{>0}$
such that
$(L_{a})_{*}\mu=j(a)\mu,\qquad(R_{a})_{*}\mu=\rho(a)\mu$
for all $a\in Q$.
Again, the right quasi-invariance condition is equivalent to
$\int_{Q}f(xa)\,d\mu(x)=\rho(a)\int_{Q}f(x)\,d\mu(x)$
for all suitable $f$.
Remark 5.
The assumption of two-sided quasi-invariance is admittedly strong. It is introduced here in order to make the cocycle calculation transparent when both left and right translations appear in the Moufang identity. The existence and naturality of such measures remain open problems.
Remark 6.
This two-sided hypothesis is strong, and we do not claim here that it holds for all locally compact topological quasigroups. Its role is methodological: it provides a clean framework in which one can derive cocycle relations from translation identities. The existence problem for such measures is discussed later as part of the research program rather than as an established theorem.
4.3 Elementary consequences
We record a simple but useful functorial fact about push-forwards.
Lemma 1.
Let $X$ be a locally compact Hausdorff space, let $\mu$ be a regular Borel measure on $X$, and let $S,T:X\to X$ be homeomorphisms. Then
$(S\circ T)_{*}\mu=S_{*}(T_{*}\mu).$
Proof.
For every Borel set $E\subseteq X$,
$((S\circ T)_{*}\mu)(E)=\mu((S\circ T)^{-1}(E))=\mu(T^{-1}(S^{-1}(E)))=(T_{*}\mu)(S^{-1}(E))=(S_{*}(T_{*}\mu))(E).$
Applying this to left and right translations gives the basic computational rule used throughout the paper.
5 The Moufang Identity and Translation Calculus
We now turn to the Moufang-type identity
$((xy)z)y=x(y(zy)).$
The key point is that this identity can be rewritten as an operator identity on $Q$.
6 Deriving Cocycle Relations from $(N1)$
We now explain in detail how $(N1)$ constrains the modular cocycles, under the additional assumption of two-sided quasi-invariance.
The Moufang identity can be expressed naturally in terms of translation operators. This formulation will be useful for the measure-theoretic computations below.
Proposition 1.
Let $Q$ be a locally compact Hausdorff topological quasigroup satisfying $(N1)$. Assume that $\mu$ is a two-sided quasi-invariant measure on $Q$, so that
$(L_{a})_{*}\mu=j(a)\mu,\qquad(R_{a})_{*}\mu=\rho(a)\mu$
for positive functions $j,\rho:Q\to\mathbb{R}_{>0}$. Then for all $x,y\in Q$,
$j(xy)\rho(y)=j(x)j(y)\rho(y).$
Since $\rho(y)>0$, this simplifies to
$j(xy)=j(x)j(y).$
Proof.
Fix $x,y\in Q$. By the operator form of $(N1)$ established above, we have
$R_{y}\circ L_{xy}=L_{x}\circ L_{y}\circ R_{y}.$
Taking push-forwards and using the functoriality lemma,
$(R_{y}\circ L_{xy})_{*}\mu=(R_{y})_{*}((L_{xy})_{*}\mu),$
and similarly
$(L_{x}\circ L_{y}\circ R_{y})_{*}\mu=(L_{x})_{*}((L_{y})_{*}((R_{y})_{*}\mu)).$
We now compute each side separately.
For the left-hand side,
$(L_{xy})_{*}\mu=j(xy)\mu$
by definition of $j$. Applying $(R_{y})_{*}$ gives
$(R_{y})_{*}((L_{xy})_{*}\mu)=(R_{y})_{*}(j(xy)\mu)=j(xy)(R_{y})_{*}\mu=j(xy)\rho(y)\mu.$
For the right-hand side, first use
$(R_{y})_{*}\mu=\rho(y)\mu.$
Then apply $(L_{y})_{*}$:
$(L_{y})_{*}((R_{y})_{*}\mu)=(L_{y})_{*}(\rho(y)\mu)=\rho(y)(L_{y})_{*}\mu=\rho(y)j(y)\mu.$
Now apply $(L_{x})_{*}$:
$(L_{x})_{*}((L_{y})_{*}((R_{y})_{*}\mu))=(L_{x})_{*}(\rho(y)j(y)\mu)=\rho(y)j(y)(L_{x})_{*}\mu=\rho(y)j(y)j(x)\mu.$
Since the two composite transformations are equal, their push-forwards of $\mu$ are equal. Hence
$j(xy)\rho(y)\mu=\rho(y)j(y)j(x)\mu.$
Because $\mu$ is nonzero and $\rho(y)>0$, it follows that
$j(xy)=j(x)j(y).$
Remark 7.
This is the cleanest rigorous form of the cocycle calculation. The right quasi-invariance hypothesis is essential for the derivation as written, because $(N1)$ mixes left and right translations. Without some control on $(R_{y})_{*}\mu$, one cannot simply cancel the right-translation term.
The proposition shows that, under two-sided quasi-invariance, the left modular cocycle behaves multiplicatively with respect to the quasigroup product. This is formally analogous to the multiplicativity of the modular function on a locally compact group.
Remark 8.
At this point one should be cautious. The identity
$j(xy)=j(x)j(y)$
does not by itself prove that $j\equiv 1$. Additional structural input is required. The main point of the present paper is that $(N1)$ naturally drives the measure-theoretic structure toward a group-like modular behavior, and that this behavior is plausibly related to loop structure.
Remark 9.
The multiplicativity relation for $j$ does not by itself imply that the cocycle is trivial. In the group case, nontrivial characters frequently exist. The possibility that Moufang-type identities impose additional rigidity on such multiplicative functions remains an interesting open question.
7 Representation of the translation group
Theorem 2 (Translation representation).
Let $Q$ be a quasigroup and let
$\chi:Q\to\mathbb{R}_{>0}$
be a multiplicative function satisfying
$\chi(xy)=\chi(x)\chi(y).$
Then the map
$\pi:\mathrm{LMlt}(Q)\to\mathbb{R}_{>0}$
defined by
$\pi(L_{a})=\chi(a)$
extends uniquely to a group homomorphism
$\pi:\mathrm{LMlt}(Q)\to\mathbb{R}_{>0}.$
Proof.
Let $\mathrm{LMlt}(Q)$ be the group generated by the left translations $L_{a}$ $(a\in Q)$.
Define $\pi$ on generators by
$\pi(L_{a})=\chi(a).$
Consider two generators $L_{a},L_{b}$. Their composition satisfies
$(L_{a}\circ L_{b})(x)=a(bx).$
Using multiplicativity of $\chi$, we compute
$\chi(a(bx))=\chi(a)\chi(bx)=\chi(a)\chi(b)\chi(x).$
Thus
$\chi((L_{a}\circ L_{b})(x))=\pi(L_{a})\pi(L_{b})\chi(x).$
Therefore the assignment
$L_{a}\mapsto\chi(a)$
is multiplicative with respect to composition of left translations.
Since $\mathrm{LMlt}(Q)$ is generated by the $L_{a}$, this determines a unique group homomorphism
$\pi:\mathrm{LMlt}(Q)\to\mathbb{R}_{>0}.$
Remark 10.
If the modular cocycle arising from a Haar-type measure is multiplicative, the theorem shows that it determines a one-dimensional representation of the left multiplication group of the quasigroup.
8 Group Case
Proposition 2 (Group case).
Let $G$ be a locally compact group and let $\mu$ be a left Haar measure. Then for every $a\in G$ there exists a positive number $\Delta(a)$ such that
$(R_{a})_{*}\mu=\Delta(a)\mu.$
The function
$\Delta:G\to\mathbb{R}_{>0}$
is a continuous group homomorphism satisfying
$\Delta(ab)=\Delta(a)\Delta(b).$
Proof.
Let $\mu$ be a left Haar measure on $G$. By left invariance we have
$(L_{a})_{*}\mu=\mu\qquad(a\in G).$
For right translations, the measure $(R_{a})_{*}\mu$ is again a left invariant measure. Since Haar measure is unique up to a positive scalar, there exists $\Delta(a)>0$ such that
$(R_{a})_{*}\mu=\Delta(a)\mu.$
Let $a,b\in G$. Using functoriality of pushforward we compute
$(R_{ab})_{*}\mu=(R_{b}\circ R_{a})_{*}\mu=(R_{b})_{*}((R_{a})_{*}\mu).$
Substituting the definition of $\Delta$ gives
$(R_{b})_{*}(\Delta(a)\mu)=\Delta(a)(R_{b})_{*}\mu=\Delta(a)\Delta(b)\mu.$
Thus
$\Delta(ab)=\Delta(a)\Delta(b).$
Continuity of $\Delta$ follows from the continuity of the translation action. ∎
Remark 11.
In the group case the modular function $\Delta$ measures the defect of right invariance of the Haar measure. The multiplicativity of $\Delta$ is a direct consequence of the associativity of the group operation.
In the quasigroup setting considered in this paper, multiplicativity of the modular cocycle is instead derived from the Moufang identity.
9 The $ax+b$ Group as a Concrete Model
We now record a standard example in full detail.
Let
$G=\{(a,b)\mid a>0,\ b\in\mathbb{R}\}$
with multiplication
$(a,b)(a^{\prime},b^{\prime})=(aa^{\prime},\,b+ab^{\prime}).$
This is the group of orientation-preserving affine transformations of the real line,
$x\mapsto ax+b.$
9.1 Left Haar measure
We claim that
$d\mu(a,b)=\frac{da\,db}{a^{2}}$
is a left Haar measure on $G$.
Fix $(\alpha,\beta)\in G$. Left translation is
$L_{(\alpha,\beta)}(a,b)=(\alpha a,\ \beta+\alpha b).$
Set
$u=\alpha a,\qquad v=\beta+\alpha b.$
Then
$a=\frac{u}{\alpha},\qquad b=\frac{v-\beta}{\alpha},$
and the Jacobian determinant is
$\left|\frac{\partial(u,v)}{\partial(a,b)}\right|=\alpha^{2}.$
Hence
$du\,dv=\alpha^{2}\,da\,db.$
Also,
$u^{2}=(\alpha a)^{2}=\alpha^{2}a^{2}.$
Therefore
$\frac{du\,dv}{u^{2}}=\frac{\alpha^{2}\,da\,db}{\alpha^{2}a^{2}}=\frac{da\,db}{a^{2}}.$
Thus the density $a^{-2}\,da\,db$ is invariant under left translation, proving that it is a left Haar measure.
9.2 Right translations and the modular function
Now fix $(\alpha,\beta)\in G$ and consider right translation:
$R_{(\alpha,\beta)}(a,b)=(a\alpha,\ b+a\beta).$
Set
$u=a\alpha,\qquad v=b+a\beta.$
Then
$a=\frac{u}{\alpha},\qquad b=v-\frac{u}{\alpha}\beta.$
The Jacobian matrix is
$\begin{pmatrix}\alpha&0\\ \beta&1\end{pmatrix}$,
so
$\left|\frac{\partial(u,v)}{\partial(a,b)}\right|=\alpha.$
Hence
$du\,dv=\alpha\,da\,db.$
Using again $u=\alpha a$, we obtain
$\frac{du\,dv}{u^{2}}=\frac{\alpha\,da\,db}{\alpha^{2}a^{2}}=\frac{1}{\alpha}\frac{da\,db}{a^{2}}.$
Equivalently,
$\frac{da\,db}{a^{2}}=\alpha\frac{du\,dv}{u^{2}}.$
Thus under right translation,
$(R_{(\alpha,\beta)})_{*}\mu=\alpha\,\mu.$
Comparing with the standard defining relation
$(R_{g})_{*}\mu=\Delta(g^{-1})\mu,$
we find
$\Delta((\alpha,\beta)^{-1})=\alpha.$
Since
$(\alpha,\beta)^{-1}=\left(\alpha^{-1},-\alpha^{-1}\beta\right),$
it follows that
$\Delta(a,b)=a^{-1}.$
This example is useful because it makes the modular defect entirely explicit. In the quasigroup setting, the cocycle $j$ should be regarded as an analogue of this phenomenon.
10 Existence Questions and the Role of the Translation Group
The preceding sections explain how a modular cocycle behaves once a suitable measure is given. The much harder question is whether such measures exist.
For groups, existence comes from the Haar theorem. For quasigroups, the lack of associativity obstructs a direct analogue. A more promising approach is to look not first at the quasigroup itself, but at the transformation groups generated by its translations.
If $\mathrm{LMlt}(Q)$ or $\mathrm{Mlt}(Q)$ carries a locally compact group topology acting continuously on $Q$, and if $Q$ can be modeled as a homogeneous space for that action, then one may hope to induce a quasi-invariant measure on $Q$ from a Haar measure on the acting group. This line of thought suggests that the correct general setting may involve transformation groups, groupoids, or measure classes glued from local translation data.
We do not attempt to solve this existence problem here. Instead, the present paper isolates the cocycle mechanism that would become relevant once such a measure is available.
11 Toward a Measure-Theoretic Reading of Kunen’s Theorem
We can now formulate the main conceptual point of the paper.
Conjecture 1.
Let $Q$ be a locally compact Hausdorff topological quasigroup satisfying $(N1)$. Suppose that $Q$ admits a nonzero two-sided quasi-invariant measure with modular cocycles $j$ and $\rho$. Then the identity $(N1)$ forces the left cocycle $j$ to satisfy strong rigidity conditions, and under suitable additional assumptions this rigidity collapses $j$ to the trivial cocycle:
$j(a)=1\qquad(a\in Q).$
If such a collapse can be established, then the quasigroup would become, from the measure-theoretic point of view, analogous to a unimodular object. The guiding philosophy is that this unimodularity should correspond to the emergence of loop structure.
Remark 12.
At present this is a research program rather than a completed proof of Kunen’s theorem. The point of the present note is to make precise the modular-cocycle framework and to show that the identity $(N1)$ interacts with it in a nontrivial and structurally meaningful way.
The purpose of this paper is to propose a framework in which Haar-type structures can be studied in the quasigroup setting using quasi-invariant measures and modular cocycles.
This paper proposes a measure-theoretic viewpoint on quasigroup structure, suggesting that loop identities may be interpreted as a form of unimodularity in the translation geometry of topological quasigroups.
12 Translation–cocycle compatibility
Theorem 3 (Translation–cocycle compatibility).
Let $Q$ be a locally compact Hausdorff topological quasigroup satisfying the Moufang identity
$((xy)z)y=x(y(zy)).$
Assume that $\mu$ is a nonzero regular Borel measure on $Q$ such that
$(L_{a})_{*}\mu=j(a)\mu,\qquad(R_{a})_{*}\mu=\rho(a)\mu$
for positive functions
$j,\rho:Q\to\mathbb{R}_{>0}.$
Then, for all $x,y\in Q$,
$j(xy)\rho(y)=j(x)j(y)\rho(y).$
Since $\rho(y)>0$, it follows that
$j(xy)=j(x)j(y).$
Proof.
By Theorem LABEL:thm:translation-form-N1, the Moufang identity $(N1)$ is equivalent to the operator identity
$R_{y}\circ L_{xy}=L_{x}\circ L_{y}\circ R_{y}\qquad(x,y\in Q).$
Fix $x,y\in Q$. Applying push-forward to both sides of this identity, and using the functoriality of push-forward, we obtain
$(R_{y}\circ L_{xy})_{*}\mu=(R_{y})_{*}((L_{xy})_{*}\mu)$
and
$(L_{x}\circ L_{y}\circ R_{y})_{*}\mu=(L_{x})_{*}((L_{y})_{*}((R_{y})_{*}\mu)).$
We now compute each side separately.
Left-hand side. Since
$(L_{xy})_{*}\mu=j(xy)\mu,$
we have
$(R_{y})_{*}((L_{xy})_{*}\mu)=(R_{y})_{*}(j(xy)\mu).$
Because $j(xy)$ is a scalar, it factors out of the push-forward:
$(R_{y})_{*}(j(xy)\mu)=j(xy)(R_{y})_{*}\mu.$
Using the right quasi-invariance relation
$(R_{y})_{*}\mu=\rho(y)\mu,$
we get
$(R_{y})_{*}((L_{xy})_{*}\mu)=j(xy)\rho(y)\mu.$
Right-hand side. First use
$(R_{y})_{*}\mu=\rho(y)\mu.$
Applying $(L_{y})_{*}$ gives
$(L_{y})_{*}((R_{y})_{*}\mu)=(L_{y})_{*}(\rho(y)\mu)=\rho(y)(L_{y})_{*}\mu=\rho(y)j(y)\mu.$
Now apply $(L_{x})_{*}$:
$(L_{x})_{*}((L_{y})_{*}((R_{y})_{*}\mu))=(L_{x})_{*}(\rho(y)j(y)\mu).$
Again the scalar factors out:
$(L_{x})_{*}(\rho(y)j(y)\mu)=\rho(y)j(y)(L_{x})_{*}\mu.$
Using
$(L_{x})_{*}\mu=j(x)\mu,$
we obtain
$(L_{x})_{*}((L_{y})_{*}((R_{y})_{*}\mu))=\rho(y)j(y)j(x)\mu.$
Since
$R_{y}\circ L_{xy}=L_{x}\circ L_{y}\circ R_{y},$
their push-forwards of $\mu$ must coincide. Therefore
$j(xy)\rho(y)\mu=\rho(y)j(y)j(x)\mu.$
Because $\mu$ is nonzero and $\rho(y)>0$, we conclude that
$j(xy)=j(x)j(y).$
This proves the theorem. ∎
13 Lemma for Normalization of multiplicative cocycles
Remark 13.
The normalization condition $\chi(e)=1$ is the quasigroup analogue of the standard normalization for characters on groups. In the present context it shows that once loop structure emerges, the modular cocycles are automatically normalized at the identity.
Lemma 2 (Normalization of multiplicative cocycles).
Let $Q$ be a quasigroup and let
$\chi:Q\to\mathbb{R}_{>0}$
be a multiplicative function, i.e.
$\chi(xy)=\chi(x)\chi(y)\qquad(x,y\in Q).$
If $Q$ has a two-sided identity element $e$, then
$\chi(e)=1.$
In particular, if a modular cocycle $j$ or $\rho$ is multiplicative on a loop, then it is normalized at the identity.
Proof.
Let $e$ be a two-sided identity element of $Q$. Then
$ee=e.$
Applying multiplicativity of $\chi$, we obtain
$\chi(e)=\chi(ee)=\chi(e)\chi(e)=\chi(e)^{2}.$
Since $\chi(e)>0$, it follows that
$\chi(e)=1.$
14 Unsolved Problems
Problem 1 (Modular cocycle rigidity).
Let $Q$ be a locally compact topological Moufang quasigroup. Suppose there exists a quasi-invariant measure $\mu$ such that
$(L_{a})_{*}\mu=j(a)\mu.$
Assume that the modular cocycle $j$ satisfies
$j(xy)=j(x)j(y).$
Under what conditions must $j$ be trivial, i.e.
$j\equiv 1?$
Problem 2.
Let $Q$ be a locally compact topological quasigroup. Under what conditions does there exist a quasi-invariant measure satisfying
$(L_{a})_{*}\mu=j(a)\mu?$
15 Existence Issues and the Unimodularity Perspective
15.1 The existence problem
In contrast to the classical Haar theorem for locally compact groups, the existence of quasi-invariant measures on general locally compact topological quasigroups is not currently known in full generality.
Accordingly, the results of this paper should be interpreted as conditional statements describing structural consequences of the existence of such measures. More precisely, throughout the paper we assume the existence of a regular Borel measure $\mu$ satisfying the quasi-invariance relations
$(L_{a})_{*}\mu=j(a)\mu,\qquad(R_{a})_{*}\mu=\rho(a)\mu.$
Under this assumption we derived a compatibility relation between the translation structure and the modular cocycle, and showed that the Moufang identity forces the multiplicativity relation
$j(xy)=j(x)j(y).$
Thus the main contribution of the present work is to identify structural consequences of the existence of quasi-invariant measures on topological quasigroups.
15.2 Finite quasigroups as a basic example
Although the existence problem is open in general, finite quasigroups provide a natural class of examples in which the framework becomes trivial but consistent.
Proposition 3.
Let $Q$ be a finite quasigroup and let $\mu$ be the counting measure
$\mu(E)=|E|.$
Then $\mu$ is invariant under all left and right translations. Consequently, the associated modular cocycle is trivial,
$j\equiv 1,\qquad\rho\equiv 1.$
Proof.
Since every translation $L_{a}$ and $R_{a}$ is a bijection of the finite set $Q$, the counting measure is preserved:
$(L_{a})_{*}\mu=\mu,\qquad(R_{a})_{*}\mu=\mu.$
Thus the cocycle factors are equal to $1$. ∎
This simple example shows that the modular cocycle may be interpreted as measuring the failure of translation invariance beyond the finite case.
15.3 Unimodularity and loop structure
In the theory of locally compact groups, the modular function
$\Delta:G\to\mathbb{R}_{>0}$
measures the defect of right invariance of Haar measure. A group is called unimodular if $\Delta\equiv 1$.
The present framework suggests a similar interpretation in the quasigroup setting. The function $j:Q\to\mathbb{R}_{>0}$ obtained from quasi-invariance plays the role of a modular cocycle.
Conjecture 2.
Let $Q$ be a locally compact topological quasigroup satisfying the Moufang identity. If a quasi-invariant measure exists and the associated modular cocycle satisfies suitable regularity conditions, then the cocycle collapses to the trivial one
$j\equiv 1.$
Such a result would suggest that loop structure may be interpreted as a form of unimodularity in the translation geometry of quasigroups.
16 Conclusion
We introduced a framework for Haar-type measures on topological quasigroups, centered on quasi-invariance under translations and the associated modular cocycles. We then showed, under explicit two-sided quasi-invariance assumptions, that the Moufang-type identity $(N1)$ yields a rigorous multiplicativity relation for the left cocycle.
This places the quasigroup situation in close analogy with the classical theory of the modular function on locally compact groups. In particular, it supports the idea that loop structure may be interpreted as a kind of unimodularity in the translation geometry of a quasigroup.
Several problems remain open, most importantly the existence theory for quasi-invariant measures on topological quasigroups and the precise mechanism by which cocycle rigidity should force an identity element. These questions suggest that the interaction among quasigroup theory, topology, and harmonic analysis deserves further study.
Remark 14.
The purpose of this note is not to provide a new proof of Kunen’s theorem. Rather, we propose a measure-theoretic framework in which the Moufang identity interacts with translation geometry through modular cocycles. We hope that this viewpoint may stimulate further work connecting quasigroup theory with harmonic analysis.
Appendix A Structural Outline of the Theory
For the convenience of the reader we summarize here the logical structure of the theory developed in this paper. The framework connects quasigroup structure, translation operators, and measure-theoretic cocycles.
Algebraic foundation
  • Definition A.1 (Quasigroup). A quasigroup is a set $Q$ with a binary operation such that for every $a,b\in Q$ the equations
    $ax=b,\qquad ya=b$
    have unique solutions.
  • Definition A.2 (Topological quasigroup). A topological quasigroup is a quasigroup endowed with a Hausdorff topology such that multiplication and division operations are continuous.
Translation operators
  • Definition A.3 (Translations).
    $L_{a}(x)=ax,\qquad R_{a}(x)=xa.$
  • Definition A.4 (Multiplication groups).
    $\mathrm{LMlt}(Q)=\langle L_{a}\mid a\in Q\rangle,\qquad\mathrm{Mlt}(Q)=\langle L_{a},R_{a}\mid a\in Q\rangle.$
Measure-theoretic framework
  • Definition A.5 (Pushforward measure).
    For a measurable map $T:X\to X$ and a measure $\mu$,
    $(T_{*}\mu)(E)=\mu(T^{-1}(E)).$
  • Lemma A.6 (Functoriality of pushforward).
    $(S\circ T)_{*}\mu=S_{*}(T_{*}\mu).$
  • Definition A.7 (Quasi-invariant measure).
    A measure $\mu$ on $Q$ is left quasi-invariant if
    $(L_{a})_{*}\mu=j(a)\mu,$
    and right quasi-invariant if
    $(R_{a})_{*}\mu=\rho(a)\mu.$
Translation form of Moufang identity
  • Theorem A.8 (Translation form of the Moufang identity).
    The Moufang identity
    $((xy)z)y=x(y(zy))$
    is equivalent to
    $R_{y}\circ L_{xy}=L_{x}\circ L_{y}\circ R_{y}.$
Modular cocycles
  • Theorem A.9 (Cocycle multiplicativity).
    Let $\mu$ be a two-sided quasi-invariant measure satisfying
    $(L_{a})_{*}\mu=j(a)\mu,\qquad(R_{a})_{*}\mu=\rho(a)\mu.$
    If $Q$ satisfies the Moufang identity, then the modular cocycle satisfies the compatibility relation
    $j(xy)\rho(y)=j(x)j(y)\rho(y).$
    In particular, since $\rho(y)>0$, we obtain
    $j(xy)=j(x)j(y).$
Multiplicative functions
  • Definition A.10 (Multiplicative function).
    A function
    $\chi:Q\to\mathbb{R}_{>0}$
    is multiplicative if
    $\chi(xy)=\chi(x)\chi(y).$
  • Lemma A.11 (Normalization).
    If $Q$ has an identity element $e$, then
    $\chi(e)=1.$
Representation of the translation group
  • Theorem A.12 (Translation representation).
    If $\chi:Q\to\mathbb{R}_{>0}$ is multiplicative, then
    $L_{a}\mapsto\chi(a)$
    extends to a homomorphism
    $\pi:\mathrm{LMlt}(Q)\to\mathbb{R}_{>0}.$
Conceptual structure
The logical structure of the theory can be summarized schematically as
quasigroup structure
$\downarrow$
translation operators
$\downarrow$
quasi-invariant measures
$\downarrow$
modular cocycles
$\downarrow$
Moufang identity
$\downarrow$
cocycle multiplicativity
$\downarrow$
multiplicative functions
$\downarrow$
$\text{representation of the translation group}.$
This framework suggests a measure-theoretic viewpoint in which loop structure may be interpreted as a form of unimodularity in the translation geometry of quasigroups.
Appendix B Structural Diagram of the Theory
For the convenience of the reader, we summarize here the logical architecture of the paper in diagrammatic form.
[…]
The guiding interpretation is that the modular cocycle measures the defect of translation invariance, while the Moufang identity imposes a rigidity condition on this defect. In this sense, loop structure may be viewed as a form of unimodularity in the translation geometry of quasigroups.
Takao Inoué
Faculty of Informatics
Yamato University
Katayama-cho 2-5-1, Suita, Osaka, 564-0082, Japan
inoue.takao@yamato-u.ac.jp
(Personal) takaoapple@gmail.com (I prefer my personal mail)